# Learning Encoding-Decoding Direction Pairs to Unveil Concepts of Influence in Deep Vision Networks

### v1

@@ Abstract
Empirical evidence shows that deep vision networks represent concepts as directions in latent space, vectors which we call concept embeddings. For each concept, a latent factor—a scalar—indicates the degree of its presence in an input patch. For a given patch, the latent factors of multiple concepts are encoded into a compact vector representation by linearly combining concept embeddings, with the latent factors serving as coefficients. Since these embeddings enable such encoding, we refer to them as encoding directions. A latent factor can be recovered from the representation by taking the inner product with a filter, a vector which we call a decoding direction. These encoding-decoding direction pairs are not directly accessible, but recovering them unlocks significant potential to open the black-box nature of deep networks, enabling understanding, debugging, and improving deep learning models. Decoding directions help attribute meaning to latent codes, while encoding directions help assess the influence of the concept on the predictions, and both directions may assist model correction by unlearning concepts irrelevant to the network’s prediction task. Compared to previous matrix decomposition, autoencoder, and dictionary learning approaches which rely on the reconstruction of feature activations, we propose a different perspective to learn these direction pairs. We base identifying the decoding directions on directional clustering of feature activations and introduce signal vectors to estimate encoding directions under a probabilistic perspective. Unlike most other works, we also take advantage of the knowledge encoded in the weights of the network to guide our direction search. For this, we illustrate that a novel technique called Uncertainty Region Alignment can exploit this knowledge to effectively reveal interpretable directions that influence the network’s predictions. We perform a thorough and multifaceted comparative analysis to offer insights on the fidelity of direction pairs, the advantages of the method compared to other unsupervised direction learning approaches, and how the learned directions compare in relation to those learned with supervision. We find that: a) In controlled settings with synthetic data, our approach is effective in recovering the ground-truth encoding-decoding direction pairs; b) In real-world settings, the decoding directions correspond to monosemantic interpretable concepts, often scoring substantially better in interpretability metrics than other unsupervised baselines; c) In the same settings, signal vectors are faithful estimators of the concept encoding directions validated with a novel approach based on activation maximization. At the application level, we provide examples that demonstrate how the learned directions can help to a) understand global model behavior; b) explain individual sample predictions in terms of local, spatially-aware, concept contributions; and c) intervene on the network’s prediction strategy to provide either counterfactual explanations or correct erroneous model behavior.
@@ 1 Introduction
Figure 1: Linear Representation Hypothesis: A hypothesis suggesting that deep networks encode high-level concepts, such as sky or boat, in distinct directions of their latent space, respectively ${\bm{s}}_{1}$ and ${\bm{s}}_{2}$. The illustration shows the encoding of two concept latent factors (i.e., the degree of concept presence) $\alpha^{1},\alpha^{2}$ within a patch’s representation ${\bm{x}}_{{\bm{p}}}$ by utilizing the concept embedding directions $\bm{s}$}. Additionally, it demonstrates how a filter ${\bm{w}}_{1}$ can be employed to extract one of these latent factors from the representation. The illustration omits depicting the latent space bias for brevity. We use the terms {encoding direction, concept embedding, signal direction} and the terms {filter, decoding direction} interchangeably throughout this article.
@@
The linear representation hypothesis suggests that deep neural networks encode high-level concepts in directions of their latent space (Bereska & Gavves, 2024). This hypothesis is sufficiently supported by empirical evidence, mostly by the effectiveness of linear probing. In the latter, a single linear layer can be trained on top of feature representations originating from the upper part of deep neural networks to solve semantic tasks with great success (Szegedy et al., 2014; Alain & Bengio, 2018; Zhou et al., 2018; Kim et al., 2018; Elhage et al., 2022; Nanda et al., 2023). In the concept encoding mechanism suggested by the linear representation hypothesis, the concept content of an input patch is written to its embedding as a linear combination of concept embeddings: directions in the latent space represented as vectors, each one associated with a concept. The corresponding scalar coefficient in the linear combination constitutes the concept’s latent factor that encodes the degree of concept presence in the input. Due to its utility in encoding the latent factor to a vector representation, we also refer to the concept embedding as the concept’s encoding or signal direction. Retrieving the latent factor back can be accomplished by taking the inner product between the patch embedding and a filter, another vector which we term a decoding direction. (Fig. 1) An encoding-decoding direction pair is interpretable whenever it is related to a concept that is aligned with human intuition; and is influential whenever it is related to a concept that the network consistently uses to make predictions. The latter can be quantified using methods for concept sensitivity testing, such as Kim et al. (2018) and Pfau et al. (2020).
On the one hand, since the decoding direction extracts the concept’s latent factor from the representation and the concept factor is related to the presence of the concept, it enables understanding of representations, attributing meaning to latent codes (Zhou et al., 2018; Kim et al., 2018). On the other hand, the encoding direction allows one to assess the influence of the concept on the network’s predictions (Fel et al., 2023c; Pahde et al., 2024), and both directions may be used to force the network to unlearn concepts irrelevant to its prediction task (Anders et al., 2022; Pahde et al., 2023; Dreyer et al., 2024a). Most previous approaches (Zhou et al., 2018; Kim et al., 2018; Zhang et al., 2021; Fel et al., 2023c; Doumanoglou et al., 2023; Pahde et al., 2024; Doumanoglou et al., 2024) usually focus on identifying decoding or encoding directions in isolation, limiting their applicability to specific appropriate tasks. Moreover, many of them do not explicitly make this distinction and consider using the concepts’ decoding directions in use cases where the encoding direction is a better fit. This has recently been pinpointed in the context of concept sensitivity testing and model correction in Pahde et al. (2024).
The encoding - decoding mechanism of concepts for a pre-trained network is not directly accessible. Instead, it is a latent mechanism that needs to be inferred by any means of reverse engineering. Top-down approaches (Zhou et al., 2018; Kim et al., 2018; Pahde et al., 2024) typically guess a concept of interest and subsequently verify whether the network encodes it by linear probing. However, this approach suffers from the need for concept speculation and the access to supervision via expensive annotations. In contrast, bottom-up, unsupervised approaches overcome these limitations, but face significant challenges. Previous unsupervised approaches that attempt to uncover such mechanisms include matrix decomposition (based on principal component analysis or non-negative matrix factorization)(Zhang et al., 2021; Graziani et al., 2023b), sparse autoencoders (SAEs) and dictionary learning (Bricken et al., 2023; Lim et al., 2024; Cunningham et al., 2024; Bussmann et al., 2025). These approaches rely on a decompose-then-reconstruct regime of feature activations. However, attempting to explain every tiny bit of information in the components of latent feature representations is particularly challenging. In addition, not all components may be of interest, as some of them may be noisy distractors (Haufe et al., 2014; Kindermans et al., 2017). Additionally, the reconstruction error of feature activations constitutes dark matter (Engels et al., 2025), which remains largely unexploited. Last, for effective utilization to downstream tasks of interpretability, it is of paramount importance that the identified direction pairs correspond to interpretable concepts, that is, concepts that align as much as possible with human intuition. However, alignment with human intuition should not be a strict criterion, as bottom-up approaches may also reveal counterintuitive concepts that are used by the network to make predictions, possibly exposing an unintended model behavior or a novel strategy to solve a task. SAEs attempt to uncover interpretable direction pairs by learning the directions with a sparsity objective in the units of latent concept factors. However, matrix decomposition approaches do not explicitly optimize for any interpretability criterion.
Unlike previous attempts that rely solely on reconstructing feature activations, in this work we take a different approach in learning the concept encoding-decoding direction pairs. In the proposed approach, we learn the direction pairs jointly, in an unsupervised manner. Decoding directions are identified by directional clustering of feature activations and serve as an anchor to estimate encoding directions under a probabilistic perspective. Additionally, we exploit the network’s strategy baked in its weights when making predictions to guide our direction search. Unlike SAEs that focus on sparsity within units of the latent factors when learning the directions, we emphasize sparsity in the soft binary semantic space of concepts. Our method identifies direction pairs without relying on feature reconstruction, sidestepping the dark matter issue of SAEs Engels et al. (2025) and ignoring noisy distractors in data Haufe et al. (2014); Kindermans et al. (2017), while not being limited by the relative sizes of concept and embedding spaces, which is another limitation when using SAEs. We model the decoding directions using the principles of the recently introduced unsupervised interpretable basis extraction (Doumanoglou et al., 2023)  [^1] Even though it is termed a basis, according to our definitions this method identifies decoding directions., a bottom-up approach that uncovers the directional structure of the embedding space through directional clustering of feature activations. Since concepts are often encoded as directions, this approach has the potential to discover the decoding mechanism of meaningful concepts. This method provides an explicit binary and linear classification rule for concept detection by learning the decoding direction together with an additional threshold to ascertain the presence of a concept. We term this classification rule as a concept detector due to its ability to detect the presence of a concept. Additionally, we introduce signal vectors that serve as estimators for the encoding direction of a concept under a probabilistic perspective, by extending the estimator of a single concept model (Kindermans et al., 2017) to multiple concepts. Furthermore, we also show that the alignment between the uncertainty region of the network, that is, the subspace where the network’s predictions are uncertain, with the uncertainty region of the concept detectors, that is, the subspace of ambiguous concept predictions, through a process that we call Uncertainty Region Alignment (Fig. 6) can increase the interpretability and influence of the discovered direction pairs.
@@
Our experimental analysis is multifaceted, providing insight into how our method relates to other unsupervised and even supervised direction learning approaches. Using the decoding directions, we assess the ability of our approach to identify interpretable and monosemantic concepts, while we validate the fidelity of signal vectors to the concepts’ encoding directions in the challenging real-world settings using a novel approach based on activation maximization (Olah et al., 2017; Nguyen et al., 2019; Fel et al., 2023a). The experiments cover synthetic data with known ground-truth and real-world setups with four families of state-of-the-art deep vision architectures. First, in a controlled setting, the experiments show the efficacy of the proposed approach in identifying the ground-truth concept encoding-decoding direction pairs when previous work fails. Second, in real-world setups, the learned decoding directions identify highly interpretable monosemantic concepts, often scoring significantly higher than the previous unsupervised state-of-the-art baselines in most of the interpretability metrics, while evidence from the experiment with activation maximization concludes that the encoding-decoding direction pair is faithful, with the learned signal vector being a reliable estimate for the concept’s encoding direction. Finally, we demonstrate the utility of our learned directions by i) integrating with the state-of-the-art to address global model behavior understanding, ii) introducing Concept Contribution Maps to provide detailed and spatially-aware local explanations, and iii) providing concrete steps to intervene on the model’s prediction strategy, enabling counterfactual explanations and correction of flawed model behavior. Figure 2 summarizes the applications of our method that we selected to discuss in this paper.
@@ 2.1 Direction Learning
We categorize related work into supervised and unsupervised direction learning approaches. In each category, we go through the previous methods, describing the limitations, differences, and similarities with the approach proposed here.
Supervised Concept Direction Learning Typical approaches Zhou et al. (2018); Kim et al. (2018) to concept direction learning make use of a binary linear classifier together with the annotations of a concept dataset. This classifier distinguishes representations of samples with the concept from those without and must rely on the concept content encoded within the representations, thereby requiring its weight vector (filter) to extract (decode) the concept factors from these representations. Known as Concept Activation Vectors (CAVs), these filter weights are approximations to the concept decoding directions, and they are not exact due to possible distractor-noise content in feature components (Haufe et al., 2014; Kindermans et al., 2017; Pahde et al., 2024). Regarding the encoding directions, assuming that the positive cluster contains representations of samples with the concept and the negative cluster contains those without it, Pattern-CAVs (PCAVs (Pahde et al., 2024)) estimate the encoding direction of the concept by computing the difference between the cluster means. These approaches, though, rely on linear probing, requiring the practitioner not only to provide annotations but also to speculate on the name of the concept to be identified. In contrast to that, our approach is unsupervised and bottom-up, reading the structure of the latent space, overcoming both concept speculation and the need for expensive annotations.
Unsupervised Concept Direction Learning A significant amount of related work in unsupervised direction learning is based on methods for matrix decomposition. These methods decompose the matrix of feature activations and may identify the encoding directions of concepts without the need for annotations, but with some limitations. For instance, Principal Component Analysis (PCA) (Graziani et al., 2023a) and Singular Value Decomposition (SVD) (Graziani et al., 2023b) are limited by orthogonality and cannot represent concepts that do not affect variance (Fel et al., 2023b). Likewise, Non-negative Matrix Factorization (NMF) (Zhang et al., 2021; Fel et al., 2023c) only assumes positive components and lacks latent space bias, which limits expressivity. While the transpose of both the PCA matrix and the matrix of the left singular vectors in SVD correspond to the concept decoding directions, NMF lacks a simple equivalent. Instead, at inference time, in which a test sample is examined for its concept content, NMF requires an optimization problem to be solved for each one of the samples, making the approach computationally more expensive than the calculation of an inner product. Our method overcomes all these limitations.
In Dictionary Learning (Bricken et al., 2023; Yun et al., 2023) and Sparse Autoencoders (SAEs) (Sharkey et al., 2022; Cunningham et al., 2024; Bussmann et al., 2025), the goal is to learn decoding-encoding directions by decomposing representations into latent factors and subsequently reconstructing them, and this is done while enforcing sparsity in units of the latent variables. The latter constrains these approaches to be applicable only in cases where the space of latent factors is sufficiently larger than the size of the embedding space. Contrariwise, our method uses a different principle for identifying direction pairs, independent of feature reconstruction, avoiding dealing with the dark matter problem of SAEs (Engels et al., 2025) and the noisy distractor components that may be present in the data (Haufe et al., 2014; Kindermans et al., 2017), while not being restricted by the size of the concept space in relation to the size of the embedding space. This is achieved by enforcing sparsity in the semantic, soft-binary vector space of concepts, instead of sparsity in the units of the latent variables. Since it is independent of feature reconstruction, the proposed approach uses a different way to estimate the concept encoding directions than SAEs, still coupled to the learned decoding directions and the feature activations themselves. Finally, our approach additionally considers linking the directions to their use by the model, a fact that, to our knowledge, is less explored by prior work. More details on the relation of our approach with SAEs can be found in Section A.15.
In contrast to the aforementioned techniques, the method of Doumanoglou et al. (2023; 2024) learns filter directions of linear classifiers, using a concept detector model as in Zhou et al. (2018); Kim et al. (2018), but without requiring supervision. Guided by a sparsity objective and the structure of the latent space, these classifiers map feature representations to a soft-binary concept space, essentially implementing a method for directional clustering. Since the method is guided by the directional structure of the latent space and concepts are often encoded as directions, this process often unveils decoding directions that correspond to highly interpretable, monosemantic concepts. We ground our approach on this model and additionally enhance it by removing orthogonality constraints, feature space standardization, and adding loss terms to a) sustain or improve the interpretability of the identified concepts and b) reduce the impact of distractor-noise on filter weights. Although Doumanoglou et al. (2024) proposed a technique to exploit the utilization of the directions by the network in direction search, our uncertainty region alignment approach shows a notable relative improvement over this previous approach (by up to 22.56% in the interpretability metrics), in 3 of 4 cases. Finally, our work also considers the estimation of concept encoding directions, an aspect that was not addressed in these previous works. More details on this comparison can be found in Section A.2.
@@ 2.2 Applications of Directions
Our learned direction pairs enable a series of applications, from global model understanding and detailed spatially-aware local explanations down to model intervention, which allows for counterfactual explanations and model correction. Typical previous enablers of such applications are the Concept Bottleneck Models (CBMs) (Koh et al., 2020). The principal idea behind CBMs is to create an inherently interpretable model which is comprised of a backbone, a concept bottleneck, and a linear head. This model makes class predictions in two steps. In the first, the concept bottleneck predicts the concept content of the image based on the features extracted from the backbone, while in the second, the head predicts the image class based on the previously extracted concept content. What makes this architecture interpretable is the fact that the concept vector, i.e, the output of the concept bottleneck, is trained to have each dimension aligned with an interpretable concept. More recent CBM variants (Yuksekgonul et al., 2023; Oikarinen et al., 2023) proposed techniques to turn any non-interpretable model into a CBM. For the latter, the original model’s backbone is kept frozen and the last linear head is discarded. Instead, a concept bottleneck is introduced as a projection layer that translates the embeddings of the backbone to the concept space. Finally, a last sparse linear head maps the concept vector to the output classes. To learn the concept bottleneck without annotations, Yuksekgonul et al. (2023); Oikarinen et al. (2023) leveraged CLIP Radford et al. (2021), whose visual prompt capabilities (Shtedritski et al., 2023) were later exploited by Benou & Raviv (2025) to train a concept bottleneck that is additionally spatially-aware. This last approach can localize concept content within the input image, improving upon previous methods that could only predict concepts at the image level rather than at the patch level. Finally, recently, Rao et al. (2024) leveraged SAEs (Cunningham et al., 2024) to learn a CBM for CLIP itself.
All these approaches require training the last sparse linear head with access to the model’s training dataset. Furthermore, some of them also require a dataset with concept annotations or utilize a foundation model (CLIP) to learn the concept bottleneck. Unlike all of them, our work is fundamentally different, enabling all the previous applications while remaining completely non-intrusive without the need to train any network components. Overall, in this work, we show that it is possible, at least to some extent, to get the benefits of CBMs without the need to train additional heads, access concept annotations, or precisely reconstruct features. We discover directions that correspond to concepts that the model already knows and subsequently we harness them in applications of mechanistic interpretability (Saphra & Wiegreffe, 2024; Kästner & Crook, 2024; Bereska & Gavves, 2024).

### v2

@@ Abstract
Empirical evidence shows that deep vision networks often represent concepts as directions in latent space with concept information written along directional components in the vector representation of the input. However, the mechanism to encode (write) and decode (read) concept information to and from vector representations is not directly accessible as it constitutes a latent mechanism that naturally emerges from the training process of the network. Recovering this mechanism unlocks significant potential to open the black-box nature of deep networks, enabling understanding, debugging, and improving deep learning models.
In this work, we propose an unsupervised method to recover this mechanism. For each concept, we explain that under the hypothesis of linear concept representations, this mechanism can be implemented with the help of two directions: the first facilitating encoding of concept information and the second facilitating decoding. Compared to previous matrix decomposition, autoencoder, and dictionary learning approaches which rely on the reconstruction of feature activations, we propose a different perspective to learn these encoding-decoding direction pairs. We base identifying the decoding directions on directional clustering of feature activations and introduce signal vectors to estimate encoding directions under a probabilistic perspective. Unlike most other works, we also take advantage of the network’s instructions encoded in its weights to guide our direction search. For this, we illustrate that a novel technique called Uncertainty Region Alignment can exploit these instructions to reveal the encoding-decoding mechanism of interpretable concepts that influence the network’s predictions.
Our thorough and multifaceted analysis shows that, in controlled, toy settings with synthetic data, our approach can recover the ground-truth encoding-decoding direction pairs. In real-world settings, our method effectively reveals the encoding-decoding mechanism of interpretable concepts, often scoring substantially better in interpretability metrics than other unsupervised baselines. Finally, we provide concrete applications of how the learned directions can help open the black box and understand global model behavior, explain individual sample predictions in terms of local, spatially-aware, concept contributions and intervene on the network’s prediction strategy to provide either counterfactual explanations or correct erroneous model behavior.
@@ 1 Introduction
Figure 1: Concept Encoding - Decoding under the Linear Representation Hypothesis: Deep networks encode high-level concepts, such as sky or boat, in distinct directions of their latent space, respectively ${\bm{s}}_{1}$ and ${\bm{s}}_{2}$. The illustration shows the encoding of two concept latent factors (i.e., the degree of concept presence) $\alpha^{1},\alpha^{2}$ within a patch’s representation ${\bm{x}}_{{\bm{p}}}$ by utilizing the concept embedding directions $\bm{s}$}. Additionally, it demonstrates how a filter ${\bm{w}}_{1}$ can be employed to extract one of these latent factors from the representation. The illustration omits depicting the latent space bias for brevity. We use the terms {encoding direction, concept embedding, signal direction} and the terms {filter, decoding direction} interchangeably throughout this article.
@@
The linear representation hypothesis suggests that deep neural networks encode high-level concepts in directions of their latent space (mi-review). This hypothesis is sufficiently supported by empirical evidence, mostly by the effectiveness of linear probing. In the latter, a single linear layer can be trained on top of feature representations originating from the upper part of deep neural networks to solve semantic tasks with great success (IntriguingProperties; ClassifierProbes; IBD; TCAV; Superposition; LinearRepresentationsNLP). In the concept encoding mechanism suggested by the linear representation hypothesis, the concept content of an input patch is written to its embedding as a linear combination of concept embeddings: directions in the latent space represented as vectors, each one associated with a concept. The corresponding scalar coefficient in the linear combination constitutes the concept’s latent factor that encodes the degree of concept presence in the input. Due to its utility in encoding the latent factor to a vector representation, we also refer to the concept embedding as the concept’s encoding or signal direction. Retrieving the latent factor back can be accomplished by taking the inner product between the patch embedding and a filter, another vector which we term a decoding direction. (Fig. 1) An encoding-decoding direction pair is interpretable whenever it is related to a concept that is aligned with human intuition; and is influential whenever it is related to a concept that the network consistently uses to make predictions (TCAV; RCAV).
The encoding - decoding mechanism of concepts for a pre-trained network is not directly accessible. Instead, it is a latent mechanism emerged from network training and needs to be inferred by any means of reverse engineering. Recovering this mechanism is of significant importance, as accessing it enables understanding representations (IBD; TCAV), assessing influence of concepts on the network’s predictions (TCAV; RCAV; CRAFT; PCAV), debugging neural networks (XAIImageNet), and correcting erroneous model behavior (RemovingCleverHans; RevealToRevise; spurious-features; GradientPenalization).
Previous works that identify concept directions rely either on concept annotations (IBD; TCAV; PCAV) or on a decompose-then-reconstruct regime of feature activations (NMFCAV; SVD; DictionaryNLP; SAE1; SAE2; matryoshka-saes). The former suffer from the need for concept speculation and the access to annotations while the latter, face the difficulty of overcoming unstructured noise that may be present in feature activations in the form of distractors (NeuroImaging; PatternNet) or dark matter (SAE-dark-matter).
@@
Unlike previous approaches, our method relies on neither annotations nor the reconstruction of activations, thereby sidestepping these earlier limitations. Instead, we identify decoding directions via directional clustering of feature activations and estimate encoding directions from a probabilistic perspective. Moreover, in contrast to most prior work, we guide our search for directions by exploiting the network’s instructions, i.e. the prediction strategy baked into its weights through a process we call Uncertainty Region Alignment. We conduct a thorough and multifaceted analysis that highlights the effectiveness of the approach in revealing the encoding-decoding mechanism of a) the ground-truth concepts in controlled and synthetic toy settings and b) interpretable monosemantic concepts in real-world studies of state-of-the-art deep vision architectures We also demonstrate the utility of our learned directions in obtaining global, local, and counter-factual explanations, as well as identifying and correcting flawed model behavior in a toy setting (Fig. 2).
@@ 2.1 Direction Learning
We categorize related work into supervised and unsupervised direction learning approaches. In each category, we go through the previous methods, describing the limitations, differences, and similarities with our approach.
Supervised Concept Direction Learning
Concept Activation Vectors (CAVs) (TCAV; IBD), use supervised linear classifiers trained on annotated concept datasets to separate representations of samples with and without a target concept. To solve the task, the resulting classifier weights (filters) must extract (decode) the concept factors from these representations; thus, they approximate concept decoding directions, although they can be inexact due to distractor-noise in the features (NeuroImaging; PatternNet; PCAV). To estimate concept encoding directions, PCAVs (PCAV) take the difference of means between representations of samples with and without the concept. These methods require both labeled data and prior specification of the concept name. In contrast to these probing approaches, the proposed method is unsupervised, revealing concepts by analyzing the structure of the latent space, and thereby avoiding both annotation costs and upfront concept specification.
Unsupervised Concept Direction Learning Existing unsupervised direction-learning methods that rely on matrix decomposition (PCA (PCA), SVD (SVD), NMF (NMFCAV; CRAFT)) can discover concept directions without annotations but have key limitations. PCA and SVD are constrained by orthogonality and miss concepts that do not influence variance (ConceptExtractionAndImportance). NMF, while non-negative, lacks latent space bias and, thus, expressivity, and it has no straightforward decoding directions. The proposed method addresses and overcomes all these limitations.
Dictionary Learning (DictionaryNLP; DictionaryNLP2) and Sparse Autoencoders (SAEs) (SAE3; SAE2; matryoshka-saes) identify encoding-decoding directions driven by the reconstruction of representations after decomposing them into sparse latent factors. They typically consider a latent space larger than the embedding space and may suffer from the difficulty of overcoming noise in the representations in the form of dark matter (SAE-dark-matter). Contrariwise, our method is independent of feature reconstruction, unrestricted from the relative sizes between the concept and the embedding space, and unlike these prior works, it implicitly links the learned directions to how the model actually utilizes them. Further comparative details with SAEs are provided in Section A.22.
Moving away from reconstructing representations, the methods of UIBE; CBE perform directional clustering of feature activations, a process that often unveils the decoding mechanism of interpretable, monosemantic concepts. We build on these previous approaches to estimate decoding directions and do so by removing orthogonality and feature standardization constraints, and adding new loss terms to maintain or improve concept interpretability and reduce the impact of distractor noise in filter weights. Compared to the previous attempt (CBE) of linking directions to their utilization by the network, our Uncertainty Region Alignment approach achieves up to 22.56% relative improvement in interpretability metrics in 3 out of 4 cases. We also address the estimation of concept encoding directions, which these prior works did not, with further details on this comparison provided in Section A.3.
@@ 2.2 Applications of Directions
Our learned direction pairs enable a series of applications, from global model understanding and detailed spatially-aware local explanations down to model intervention, which allows for counterfactual explanations and model correction. Typical previous enablers of such applications are the Concept Bottleneck Models (CBMs) (CBMs). Classic CBMs consist of a backbone feature extractor, a concept bottleneck predicting concept content from features, and a linear head mapping concepts to classes. What makes this architecture interpretable is the fact that the concept vector, i.e, the output of the concept bottleneck, is trained to have each dimension aligned with an interpretable concept. Recent post-hoc CBM variants (PostHocCBM; LabelFreeCBM) retrofit this structure onto existing black-box models by freezing the original backbone, discarding its head, inserting a concept bottleneck projection into concept space, and learning a task-specific sparse linear head over concepts. To learn concepts without annotations, these works exploit CLIP CLIP, and further advances spatial-cbms use visual prompting (clip-visual-prompt) to make the bottleneck spatially aware, allowing localization of concepts at the patch level rather than only at the image level. Most recently (discover-then-name-sae-cbms), SAEs (SAE2) have been used to build CBMs directly for CLIP.
To access the advantages of CBMs, these previous approaches require training the last linear head using the model’s training data and, sometimes, additional concept-annotated datasets or external models like CLIP. In contrast, the proposed work is non-intrusive: it needs no training of new model components and no concept labels. It shows that one can capture many benefits of concept bottleneck models by identifying directions corresponding to concepts the model already encodes and then using these directions in applications of mechanistic interpretability (mechanistic; mechanistic-interpretability-1; mi-review).